If mr.M has a cup of water with room temperature of 20 °C but adds 80 g of ice to the drink, so it cools to 6°C and adds more ice to drink of water so the temperature is 10°C what will the volume of the drink be?
1
Expert's answer
2019-11-22T11:14:51-0500

1.

"cM(T_w-T_1)=cm(T_1-T_i)+mL"

"M(4200)(20-6)=80(4200(6-0)+334000)"

Initial mass of water:

"M=489\ g"

2.


"c(M+m)(T_2-T_1)=cm'(T_w-T_2)"

"(M+m)(T_2-T_1)=m'(T_w-T_2)"

"(489+80)(10-6)=m'(20-10)"

"m'=228\ g"

Total mass:


"M+m+m'=489+80+228=797\ g"

Total volume is


"v=0.797\approx0.8\ L"